# Hu Hosseini-Asl Chen Casanova Ghosh Żelasko Chen Li Balam Ginsburg NVIDIAUSA

### v3

@@ 4.1 Single-turn synthetic and real spoken QA
[…]
@@ 4.3 Conversational data
To enhance the model’s conversation ability on daily topics, we create Topic and UltraChat datasets (totaling 3.3k hours as shown in Table 1). For both datasets, we first generate 4-turn text-based conversations and then synthesize them using a TTS model [26]. For Topic, we randomly choose a topic between user and agent and prompt the Meta-Llama-3.1-70B-Instruct model [4] to generate a conversation. The topics are randomly chosen from the everyday-conversation dataset [30], which covers 63 everyday and science topics. To generate concise replies for efficient training, we restrict the words of each turn to be 30 words in the prompt. The generated conversations are then synthesized into speech and prepared to the duplex data format. For UltraChat, we randomly sample a chat conversation from the UltraChat dataset [31] to use as contextual information in the prompt to generate a 4-turn conversation similar to Topic.
@@ 5.2 Evaluation Data and Metrics
We evaluate turn-taking ability and speech generation quality using the UltraChat and Impatient datasets. We use two types of metrics to measure the turn-taking ability: barge-in performance and 1st response latency (see Table 2). For barge-in performance, we introduce the following metrics: 1) Barge-in latency: The time delay between the user’s speech onset and the agent stopping its response; 2) Success rate: The percentage of cases where the agent successfully stops speaking within 1.5 seconds after the user interruption; and 3) False alarm rate: The frequency at which the agent incorrectly barges in while the user speaks. Additionally, if the user stops speaking within 0.1s, the event is not counted as a false alarm, as we found that Moshi tends to proactively respond. The 1st response latency is defined as the time taken by the agent to respond to the 1st user turn. To evaluate the speech quality, we compute the UTMOS [33] using the generated agent speech after removing silence.
@@ 6.1 Conversation and Speech Generation Quality
[…]

### v4

@@ 4.1 Single-turn synthetic and real spoken QA
We create duplex training data from the aforementioned user-agent QA pairs. First, we split a pair of utterances into two streams, corresponding to the user and agent portions separately, and then insert silence into the agent stream when the user speaks, and vice versa. This gives us two streams of speech (shown as the first turn in Fig.2). This duplex structure enables the model to listen and speak simultaneously at any time. To prevent the agent from barging in unexpectedly, we insert a 0.64s silence between user and agent before the agent speaks. This duration is chosen to balance modeling of user pauses with low turn-taking latency.
@@ 4.3 Conversational data
To enhance the model's conversation ability on daily topics, we create Topic and UltraChat datasets (totaling 3.3k hours as shown in Table 1). For both datasets, we first generate 4-turn text-based conversations and then synthesize them using a TTS model [26]. For Topic, we randomly choose a topic between user and agent and prompt the Meta-Llama-3.1-70B-Instruct model [4] to generate a conversation. The topics are randomly chosen from the everyday-conversation dataset [30], which covers 63 everyday and science topics. To generate concise replies for efficient training, we restrict the words of each turn to be 30 words in the prompt. The generated conversations are then synthesized into speech and prepared to the duplex data format. For UltraChat, we randomly sample a chat conversation from the UltraChat dataset [31] to use as contextual information in the prompt to generate a 4-turn conversation similar to Topic.
@@ 5.2 Evaluation Data and Metrics
We evaluate turn-taking ability and speech generation quality using the UltraChat and Impatient datasets. We use two types of metrics to measure the turn-taking ability: barge-in performance and 1st response latency (see Table 2). For barge-in performance, we introduce the following metrics: 1) Barge-in latency: The time delay between the user's speech onset and the agent stopping its response; 2) Success rate: The percentage of cases where the agent successfully stops speaking within 1.5 seconds after the user interruption; and 3) False alarm rate: The frequency at which the agent incorrectly barges in while the user speaks. Additionally, if the user stops speaking within 0.1s, the event is not counted as a false alarm, as we found that Moshi tends to proactively respond. The 1st response latency is defined as the time taken by the agent to respond to the 1st user turn. To evaluate the speech quality, we compute the UTMOS [33] using the generated agent speech after removing silence.
@@ 6.1 Conversation and Speech Generation Quality
We cannot directly compare our 1st response latency with Moshi's as Moshi almost always responds before the user finishes talking and thus does not fit for this metric. We also note that our 1st response latency is affected by our data generation, as we always add a 0.64-second silence after the user turns to ensure no unexpected agent barge-in. Lastly, we report UTMOS and our model generates better quality speech than Moshi by up to 0.4.